w=dependent variable

v=independendt variable


w(2v-w+1)dv+v(3v-4w+3)dw=0


1
Expert's answer
2021-07-19T10:45:47-0400

"w(2v-w+1)dv+v(3v-4w+3)dw=0"


"(2vw-w^2+w)dv+(3v^2-4vw+3v)dw=0"


"M(v,w)=(2vw-w^2+w) \implies {\partial M(v,w) \over \partial w}=2v-2w+1"


"N(v,w)=(3v^2-4vw+3v) \implies {\partial N(v,w) \over \partial v}=6v-4w+3"


The differential equation is not exact since "{\partial M(v,w) \over \partial w} \not = {\partial N(v,w) \over \partial v}"

There is no integrating factor of the form "p(v)" or "p(w)"

Let's try another method

"(2vw-w^2+w)dv+(3v^2-4vw+3v)dw=0"


"(2vw-w^2+w){dv\over dv}+(3v^2-4vw+3v){dw\over dv}=0"


"(2vw-w^2+w)+(3v^2-4vw+3v){dw\over dv}=0"


"(2vw-w^2+w)+(3v^2{dw\over dv}-4vw{dw\over dv}+3v{dw\over dv})=0"


"2vw-w^2+w+3v^2{dw\over dv}-4vw{dw\over dv}+3v{dw\over dv}=0 ..........(i)"


Let "u=vw\implies {du\over dv}=w+v{dw\over dv} \implies {dw\over dv}={{du\over dv}\over v}-{u\over v^2}" and Substituting this to equation "(i)"


"2v{u\over v}-({u\over v})^2+({u\over v})+3v^2({{du\over dv}\over v}-{u\over v^2})-4v{u\over v}({{du\over dv}\over v}-{u\over v^2}) +3v({{du\over dv}\over v}-{u\over v^2})=0"


The Variables cannot be separated further

Lets check if the differential equation is homogenous


"(2vw-w^2+w)+(3v^2-4vw+3v){dw\over dv}=0"


"(3v^2-4vw+3v){dw\over dv}=-(2vw-w^2+w)"



"{dw\over dv}=-{(2vw-w^2+w)\over (3v^2-4vw+3v)}"

The differential equation is not homogenous

"\therefore" The differential equation has no solution